OpenStudy (anonymous):

Let u = <-4, -3>. Find the unit vector in the direction of u, and write your answer in component form.

OpenStudy (anonymous):

\[û = \frac{ u }{ \left| \left| u \right| \right| }\] ||u|| is the modulus of u. do you know how to find this?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

i meant to say magnitude not modulus. and the magnitude is the length of the vector we have \[u = <u_i , u_j > = <-4, -3>\]\[||u|| = \sqrt{u_i^2+ u_j^2}\] are you familiar with i and j notation?

OpenStudy (anonymous):

ummmm, no...I take this class online, I dont have a teacher to teach me this stuff

OpenStudy (anonymous):

ok. well <i, j> is usually equivalent to <x , y>. with the vector u = <-4, -3> -4 is its x component [or i component] -3 is its y component [or j] vector looks like this  ||u|| = 5

OpenStudy (anonymous):

\[û = \frac{ <-4, -3> }{ 5 } = <-\frac{ 4 }{ 5 }, -\frac{ 3 }{ 5 }>\]

OpenStudy (anonymous):

which is often written as: \[û = -\frac{ 4 }{ 5 } i - \frac{ 3 }{ 5 }j\]